OpenStudy (anonymous):

simplify this 3x3(4x2)

OpenStudy (anonymous):

Can you clarify the question, is that \(3x^3(4x^2)\)?

OpenStudy (anonymous):

yeah it is

OpenStudy (anonymous):

Okay, was confused. Just multiply it out. \(3x^3\cdot4x^2\).

OpenStudy (anonymous):

so 432?

OpenStudy (anonymous):

How did you get that? Where did the \(x\)s go?

OpenStudy (anonymous):

idk , I don't know how to do this

OpenStudy (anonymous):

432 is \(3^3\times4^2\), but you have \(3x^3\cdot4x^2\), you can't just ignore the \(x\)s.

OpenStudy (anonymous):

\(3x^3(4x^2)=3\cdot x^3\cdot4\cdot x^2=(3\cdot4)(x^3\cdot x^2)\)

OpenStudy (anonymous):

\(x^a\cdot x^b=x^{a+b}\)